# and Regional Knowledge of Kazakhstan

### v1

@@ Abstract
Mukhammed Togmanovthanks: These authors contributed equally.  Nurdaulet Mukhituly  Diana Turmakhan Jonibek MansurovMaiya GoloburdaAkhmed SakipZhuohan Xie Yuxia WangBekassyl SyzdykovAlham Fikri AjiEkaterina Kochmar Preslav NakovFajri Koto Department of Natural Language Processing, MBZUAI {mukhammed.togmanov,nurdaulet.mukhituly,diana.turmakhan}@mbzuai.ac.ae
KazMMLU: Evaluating Language Models on Kazakh, Russian,
and Regional Knowledge of Kazakhstan
Mukhammed Togmanovthanks: These authors contributed equally.  Nurdaulet Mukhituly  Diana Turmakhan Jonibek MansurovMaiya GoloburdaAkhmed SakipZhuohan Xie Yuxia WangBekassyl SyzdykovAlham Fikri AjiEkaterina Kochmar Preslav NakovFajri Koto Department of Natural Language Processing, MBZUAI {mukhammed.togmanov,nurdaulet.mukhituly,diana.turmakhan}@mbzuai.ac.ae
@@ 1 Introduction
With a population exceeding twenty million, the Republic of Kazakhstan in Central Asia remains underrepresented in the field of natural language processing (NLP) Joshi et al. (2020). This gap is highlighted by the limited progress in developing large language models (LLMs) and evaluation benchmarks specifically tailored to the languages and the cultural context of Kazakhstan. Kazakh, a Turkic language spoken by more than fourteen million people (around 70% of the population), holds substantial cultural and geopolitical significance in Central Asia. Russian, used by approximately 15% of the population, serves as the country’s second primary language. [^2] https://glottolog.org/ Although Kazakh appears in certain multilingual datasets (Yeshpanov et al., 2024, 2022; Yeshpanov and Varol, 2024), most of these resources rely heavily on translations from English, lacking the cultural richness essential for inclusive LLM development. Previous work has primarily addressed classic NLP tasks such as named entity recognition Yeshpanov et al. (2022) and sentiment analysis Yeshpanov and Varol (2024). Meanwhile, recent developments in LLM research have shifted toward more reasoning-focused evaluation Meta AI (2024); OpenAI (2024), highlighting a clear research gap for inclusive NLP in the Kazakh context. Here, we aim to bridge this gap. In particular, we introduce KazMMLU, a curated dataset of school- and university-level questions from Kazakhstan, available in both Kazakh and Russian. KazMMLU follows the framework of the Massive Multitask Language Understanding (MMLU) dataset Hendrycks et al. (2021); Koto et al. (2024); Li et al. (2024); Koto et al. (2023), which features multiple-choice questions across various subjects and education levels. MMLU has become a standard benchmark for evaluating LLMs’ reasoning and knowledge capabilities Meta AI (2024); Team et al. (2024); Qwen et al. (2025). Unlike general MMLU, KazMMLU incorporates Kazakhstan-specific content, including topics on Kazakh history, traditions, and linguistics, while also reflecting the country’s multilingual landscape by providing questions in both Kazakh and Russian. As shown in Figure 1, the dataset is divided into two categories: High School and University. KazMMLU consists of approximately 48% of the questions in Kazakh and 52% in Russian. The High School section includes questions in both Kazakh and Russian, covering subjects such as Mathematics, Physics, and Kazakh Literature. The University section only features questions in Russian, focusing on professional disciplines such as Law, Economics, and Medicine. This structure aligns with Kazakhstan’s bilingual education system and provides a more representative benchmark for evaluating LLMs in the region. KazMMLU is sourced from authentic educational materials, including national exams, textbooks, and professional certification repositories. Each question is accompanied by metadata, including the subject, level, source, and correct answer key, ensuring transparency and usability for downstream evaluations. Our contributions can be summarized as follows:
@@
  • We evaluate various multilingual LLMs, including Llama-3.1 Meta AI (2024), Qwen Qwen et al. (2025), GPT-4o OpenAI (2024), BLOOMZ Muennighoff et al. (2023), mT0 Muennighoff et al. (2023), and DeepSeek V3 DeepSeek-AI (2024), across different model sizes.
@@ 2 Related Work
Prominent models such as OpenAI’s ChatGPT, Anthropic’s Claude, and Yandex’s Yandex-GPT are designed to handle multiple languages, including Russian and Kazakh, enabling a wide range of applications from translation to content generation (OpenAI, 2024; Anthropic, 2023; Yandex, 2023). Additionally, open-source models like Meta’s Llama series provide multilingual support (Meta AI, 2024). While these models can produce text in Kazakh, they were not specifically trained or fine-tuned for it. In contrast, the Aya model, an open-access multilingual LLM, supports 101 languages, including Kazakh (Üstün et al., 2024). Kazakh-specific language models have been scarce, with most multilingual models offering only limited support. To bridge this gap, ISSAI (2024) introduced KazakhLLM, a model fine-tuned on Kazakh data from Llama, though its evaluation has primarily relied on machine-translated datasets.
@@
Evaluating LLMs across diverse linguistic and cultural contexts is increasingly critical; however, existing benchmarks overlook Kazakhstan. While benchmarks such as XCOPA (Hu et al., 2021), XFGLUE (Liang et al., 2020), and XTREME (Hu et al., 2020) assess cross-lingual performance, they exclude Kazakh, and GlobalMMLU (Singh et al., 2024) lacks Kazakhstan-specific content. In contrast, several Kazakh-specific datasets exist, including KazNERD (Yeshpanov et al., 2022) for named entity recognition, KazQAD (Yeshpanov et al., 2024) for question answering, and KazSANDRA (Yeshpanov and Varol, 2024) for sentiment analysis. However, these datasets focus on narrow tasks and do not assess reasoning, factual recall, or domain-specific knowledge. To address these limitations, KazMMLU presents a large-scale, Kazakhstan-specific benchmark covering STEM, humanities, and social sciences. Unlike previous datasets, KazMMLU supports a holistic evaluation of reasoning and domain-specific knowledge, offering a more accurate assessment of multilingual LLM capabilities and advancing AI for low-resource languages. To further illustrate the differences between KazMMLU and previous benchmarks, we compare it with two existing datasets, SIGTURK (Maxutov et al., 2024) and INCLUDE (Romanou et al., 2024), in Table 1. As shown, KazMMLU is the only dataset that incorporates real-world educational materials, professional subjects, and domain-specific reasoning in both Kazakh and Russian, offering a more localized and comprehensive evaluation of LLMs. This structured assessment underscores the importance of country-specific benchmarks in multilingual NLP research and contributes to bridging the gap in Kazakh language understanding.
@@
Model Benchmarking 41 LLMs (GPT-4o, Llama-3.1, DeepSeek V3, etc.) 7 models on Kazakh NLP tasks Multiple LLMs across 44 languages
@@ 3 KazMMLU
KazMMLU covers a wide range of subjects spanning multiple disciplines, including STEM, humanities, social sciences, and professional studies. Figure 2 illustrates the distribution of questions across different subjects. The dataset exhibits a strong representation in STEM fields, with subjects like Biology, Mathematics and Physics accounting for a significant portion of the total questions. Humanities and social sciences are also well represented, particularly Kazakh History, World History, and Law, reflecting the importance of these disciplines in the Kazakh educational system. Additionally, KazMMLU includes questions from professional domains such as Economics, Finance, Jurisprudence, and Medicine, enabling the evaluation of language models in specialized areas. The inclusion of both Kazakh and Russian language subjects further ensures a balanced linguistic representation. This diverse subject distribution provides a robust benchmark for evaluating multilingual language models across various domains. To illustrate the question format in KazMMLU, Figure 3 presents a sample multiple-choice question in Kazakh. Answering this question requires an understanding of Kazakhstan’s local context, as it covers topics such as history and geography. This example highlights the dataset’s diversity across subjects and difficulty levels.
Figure 3: Examples of a Kazakh history question and a geography test question from KazMMLU. The left side presents the original text, while the right side provides the English translation for reference. The bolded options indicate the correct answers.
@@ 3.1 Data Construction
To construct KazMMLU, we adopted a systematic approach inspired by MMLU datasets (Koto et al., 2023, 2024; Li et al., 2024). The dataset comprises questions sourced from national exams, textbooks, and professional certification materials such as iTest.kz, ymnik.kz, oltest.kz and Book - Shyn Kitap. To ensure diversity, we employed three data collection strategies: (1) automated online crawling, (2) manual transcription from scanned books, and (3) manual extraction from online sources. A detailed breakdown of dataset sources is provided in Appendix 6. For automatic online crawling, we collect question texts, multiple-choice options, correct answer keys, and metadata. For books, authors manually scan materials and apply document processing for machine-readable conversion. Two expert workers fluent in Kazakh and Russian manually extract questions from scanned books and online sources, recording metadata such as source, country, subject, level, and answer key. We compiled 23,000 questions, with (85%) from automated crawling and the rest from manual extraction. Only questions with valid answer keys were included, while multimodal ones requiring images or videos were excluded. For context-dependent questions, annotators ensured necessary context was included. A training workshop clarified guidelines, and weekly check-ins ensured consistency. Annotators were competitively compensated to maintain quality.
@@ 3.2 Quality Control
Our quality control process primarily targets the automatically-crawled data, as the other two data collection strategies involve direct human involvement. To ensure accuracy, we recruited two professional annotators, each holding at least a bachelor’s degree and fluent in both Kazakh and Russian. They manually reviewed all questions to verify correctness and completeness. Any question containing errors or missing components (e.g., incomplete contexts or broken answer options) was discarded. Through this extensive human verification, every question included in KazMMLU undergoes manual validation, ensuring a high-quality dataset. Additionally, we developed scripts to detect duplicates, verify metadata, and eliminate errors like duplicate questions, incorrect answer keys, and formatting issues, enhancing dataset reliability
@@ 3.3 Data Statistics
KazMMLU comprises 23,000 multiple-choice questions spanning two educational levels: high school and university. As shown in Figure 1, 48% of the dataset consists of high school-level questions in Kazakh, 36% in Russian, and 16% university-level questions in Russian. Table 2 outlines the subject distribution, covering STEM, Humanities, Social Sciences, and Languages. Notably, the Humanities, Social Sciences, and Language sections contain extensive Kazakhstan-specific knowledge. The dataset maintains a balanced distribution between Kazakh and Russian, reflecting the bilingual nature of Kazakhstan’s education system. Table 3 presents the average question and answer lengths across educational levels and subject areas. While the overall question length remains relatively consistent between high school and university levels, answer lengths (in characters) tend to be longer at the university level. Additionally, questions in Humanities and STEM subjects are generally longer compared to those in Social Sciences and Languages.
@@ 4.1 Setup
We evaluate 26 multilingual LLMs of various sizes in both zero-shot and few-shot settings. Our selection includes models from diverse architectures, such as BLOOM Scao et al. (2023), BLOOMZ, and mT0 Muennighoff et al. (2023), Falcon Almazrouei et al. (2023), Llama-3.1 Touvron et al. (2023a, b), GPT-4o OpenAI (2024), mT5 Xue et al. (2020), Vikhr Nikolich et al. (2024), and DeepSeek V3 DeepSeek-AI (2024). For the evaluation purpose, we use two distinct prompt configurations to examine the effect of prompt language: (1) a Kazakh prompt with English (Latin script) alphabetic output and (2) an English prompt with English alphabetic output, as illustrated in Figure 4. For placeholders such as [Subject] and [Level], we use Kazakh in the Kazakh prompt and translate them into English for the English prompt. However, the question and answer choices remain in their original language (Kazakh or Russian). A complete example prompt is provided in Appendix B (Figure 10). Following prior studies (Koto et al., 2023; Li et al., 2024), we adopt different answer selection methods based on model accessibility. For open-weight models, we apply the next-token prediction approach, computing probabilities for each multiple-choice option (A, B, C, D, or E) and selecting the one with the highest probability. This method is well-suited for autoregressive models that perform token-wise scoring. For closed-weight models (e.g., GPT–4o, Yandex-GPT, and DeepSeek V3), we use a free-text generation strategy, prompting the model to generate a textual response, from which the predicted answer is extracted via string matching. This approach is necessary due to the lack of direct token probability outputs in closed-source APIs. For evaluation, we use accuracy as the primary metric, following prior studies (Koto et al., 2023; Li et al., 2024).

### v2

@@ Abstract
Despite having a population of twenty million, Kazakhstan’s culture and language remain underrepresented in the field of natural language processing. Although large language models (LLMs) continue to advance worldwide, progress in Kazakh language has been limited, as seen in the scarcity of dedicated models and benchmark evaluations. To address this gap, we introduce KazMMLU, the first MMLU-style dataset specifically designed for Kazakh language. KazMMLU comprises 23,000 questions that cover various educational levels, including STEM, humanities, and social sciences, sourced from authentic educational materials and manually validated by native speakers and educators. The dataset includes 10,969 Kazakh questions and 12,031 Russian questions, reflecting Kazakhstan’s bilingual education system and rich local context. Our evaluation of several state-of-the-art multilingual models (Llama3.1, Qwen-2.5, GPT-4, and DeepSeek V3) demonstrates substantial room for improvement, as even the best-performing models struggle to achieve competitive performance in Kazakh and Russian. These findings highlight significant performance gaps compared to high-resource languages. We hope that our dataset will enable further research and development of Kazakh-centric LLMs [^1] KazMMLU can be accessed at https://huggingface.co/datasets/MBZUAI/KazMMLU..
@@
Mukhammed Togmanov1thanks: These authors contributed equally. Nurdaulet Mukhituly1  Diana Turmakhan1 Jonibek Mansurov1Maiya Goloburda1Akhmed Sakip1Zhuohan Xie1 Yuxia Wang1Bekassyl Syzdykov1Nurkhan Laiyk1Alham Fikri Aji1 Ekaterina Kochmar1Preslav Nakov1,2Fajri Koto1 1Mohamed bin Zayed University of Artificial Intelligence 2Institute of Foundation Models {mukhammed.togmanov,nurdaulet.mukhituly,diana.turmakhan}@mbzuai.ac.ae
@@ 1 Introduction
With a population exceeding twenty million, the Republic of Kazakhstan in Central Asia remains underrepresented in the field of natural language processing (NLP) Joshi et al. (2020). This gap is highlighted by the limited progress in developing large language models (LLMs) and evaluation benchmarks specifically tailored to the languages and the cultural context of Kazakhstan. Kazakh, a Turkic language spoken by more than fourteen million people (around 70% of the population), holds substantial cultural and geopolitical significance in Central Asia. Russian, used by approximately 15% of the population, serves as the country’s second primary language. [^2] https://glottolog.org/
Although Kazakh appears in certain multilingual datasets (Yeshpanov et al., 2024, 2022; Yeshpanov and Varol, 2024), most of these resources rely heavily on translations from English, lacking the cultural richness essential for inclusive LLM development. Previous work has primarily addressed classic NLP tasks such as named entity recognition Yeshpanov et al. (2022) and sentiment analysis Yeshpanov and Varol (2024). Meanwhile, recent developments in LLM research have shifted toward more reasoning-focused evaluation Meta AI (2024); OpenAI (2024), highlighting a clear research gap for inclusive NLP in the Kazakh context.
Here, we aim to bridge this gap. In particular, we introduce KazMMLU, a curated dataset of school- and university-level questions from Kazakhstan, available in both Kazakh and Russian. KazMMLU follows the framework of the Massive Multitask Language Understanding (MMLU) dataset Hendrycks et al. (2021); Koto et al. (2024); Li et al. (2024); Koto et al. (2023), which features multiple-choice questions across various subjects and education levels. MMLU has become a standard benchmark for evaluating LLMs’ reasoning and knowledge capabilities Meta AI (2024); Team et al. (2024); Yang et al. (2025). Unlike general MMLU, KazMMLU incorporates Kazakhstan-specific content, including topics on Kazakh history, traditions, and linguistics, while also reflecting the country’s multilingual landscape by providing questions in both Kazakh and Russian.
As shown in Figure 1, the dataset is divided into two categories: High School and University. KazMMLU consists of approximately 48% of the questions in Kazakh and 52% in Russian. The High School section includes questions in both Kazakh and Russian, covering subjects such as Mathematics, Physics, and Kazakh Literature. The University section only features questions in Russian, focusing on professional disciplines such as Law, Economics, and Medicine. This structure aligns with Kazakhstan’s bilingual education system and provides a more representative benchmark for evaluating LLMs in the region. KazMMLU is sourced from authentic educational materials, including national exams, textbooks, and professional certification repositories. Each question is accompanied by metadata, including the subject, level, source, and correct answer key, ensuring transparency and usability for downstream evaluations.
Our contributions can be summarized as follows:
@@
  • We evaluate various multilingual LLMs, including Llama3.1 Meta AI (2024), Qwen Yang et al. (2025), GPT-4o OpenAI (2024), BLOOMZ Muennighoff et al. (2023), mT0 Muennighoff et al. (2023), and DeepSeek V3 DeepSeek-AI (2024), across different model sizes.
@@ 2 Related Work
Prominent models such as OpenAI’s ChatGPT, Anthropic’s Claude, and Yandex’s Yandex-GPT are designed to handle multiple languages, including Russian and Kazakh, enabling a wide range of applications from translation to content generation (OpenAI, 2024; Anthropic, 2023; Yandex, 2023). Additionally, open-source models like Meta’s Llama series provide multilingual support (Meta AI, 2024). While these models can produce text in Kazakh, they were not specifically trained or fine-tuned for it. In contrast, the Aya model, an open-access multilingual LLM, supports 101 languages, including Kazakh (Üstün et al., 2024).
Kazakh-specific language models remain limited, with most multilingual models providing only partial support. To address this, ISSAI (2024) and Koto et al. (2025) introduced KazakhLLM and Sherkala, both fine-tuned on Kazakh data from Llama. However, their evaluations have largely relied on machine-translated benchmarks.
@@
Evaluating LLMs across diverse linguistic and cultural contexts is increasingly critical; however, existing benchmarks overlook Kazakhstan. While benchmarks such as XCOPA (Ponti et al., 2020), XGLUE (Liang et al., 2020), and XTREME (Hu et al., 2020) assess cross-lingual performance, they exclude Kazakh, and GlobalMMLU (Singh et al., 2024) lacks Kazakhstan-specific content. Recent efforts like Kardeş-NLU (Senel et al., 2024) explore cross-lingual transfer for Turkic languages including Kazakh, but focus on NLU tasks specifically, without domain-specific reasoning or local context. In contrast, several Kazakh-specific datasets exist, including KazNERD (Yeshpanov et al., 2022) for named entity recognition, KazQAD (Yeshpanov et al., 2024) for question answering, and KazSANDRA (Yeshpanov and Varol, 2024) for sentiment analysis. However, these datasets focus on narrow tasks and do not assess reasoning, factual recall, or domain-specific knowledge. TUMLU (Isbarov et al., 2025) provides a native-language benchmark for Turkic languages including Kazakh, but lacks Kazakhstan-specific context or domain-specific evaluation. To address these limitations, KazMMLU presents a large-scale, Kazakhstan-specific benchmark covering STEM, humanities, and social sciences. Unlike previous datasets, KazMMLU supports a holistic evaluation of reasoning and domain-specific knowledge, offering a more accurate assessment of multilingual LLM capabilities and advancing AI for low-resource languages.
To further illustrate the differences between KazMMLU and previous benchmarks, we compare it with two existing datasets, SIGTURK (Maxutov et al., 2024) and INCLUDE (Romanou et al., 2024), in Table 1. As shown, KazMMLU is the only dataset that incorporates real-world educational materials, professional subjects, and domain-specific reasoning in both Kazakh and Russian, offering a more localized and comprehensive evaluation of LLMs. This structured assessment underscores importance of country-specific benchmarks in multilingual NLP research and contributes to bridging the gap in Kazakh language understanding.
@@
Model Benchmarking 41 LLMs (GPT-4o, Llama3.1, DeepSeek V3, etc.) 7 models on Kazakh NLP tasks Multiple LLMs across 44 languages
@@ 3 KazMMLU
KazMMLU spans a broad range of subjects across STEM, humanities, social sciences, and professional studies. As shown in Figure 2, the dataset has strong coverage in STEM fields, with Biology, Mathematics, and Physics comprising a huge portion of the questions. Humanities and social sciences are also well represented, particularly Kazakh History, World History, and Law, reflecting their central role in the Kazakh educational system.
In addition to academic subjects, KazMMLU includes questions from professional domains such as Economics, Finance, Jurisprudence, and Medicine, enabling the evaluation of language models in specialized areas. The dataset features content in both Kazakh and Russian, ensuring balanced linguistic representation. All questions are written in Cyrillic script, which remains the prevailing standard in Kazakhstan’s formal education system. This subject and language coverage makes KazMMLU a robust benchmark for evaluating multilingual language models across diverse domains.
To illustrate the question format in KazMMLU, Figure 3 presents a sample multiple-choice question in Kazakh. Answering this question requires an understanding of Kazakhstan’s local context, as it covers topics such as history and geography. This example highlights the dataset’s diversity across subjects and difficulty levels.
Figure 3: Examples of a Kazakh History question and a Geography question from KazMMLU. The left side shows the original text, and the right side provides the English translation for reference. Correct answers are marked in bold.
@@ 3.1 Data Construction
To construct KazMMLU, we adopted a systematic approach inspired by MMLU datasets (Koto et al., 2023, 2024; Li et al., 2024). The dataset comprises questions sourced from national exams, textbooks, and professional certification materials such as iTest.kz, ymnik.kz, oltest.kz and Book - Shyn Kitap. To ensure diversity, we employed three data collection strategies: (1) automated online crawling, (2) manual transcription from scanned books, and (3) manual extraction from online sources. A detailed breakdown of dataset sources is provided in Appendix 6.
For automatic online crawling, we collect question texts, multiple-choice options, correct answer keys, and metadata. For books, authors manually scan materials and apply document processing for machine-readable conversion. Two expert workers fluent in Kazakh and Russian manually extract questions from scanned books and online sources, recording metadata such as source, country, subject, level, and answer key. In total, we compiled 23,000 questions—85% obtained through automated crawling and the remainder through manual extraction.
Only questions with valid answer keys were included, while multimodal ones requiring images or videos were excluded. For context-dependent questions, annotators ensured necessary context was included. A training workshop clarified guidelines, and weekly check-ins ensured consistency. Annotators were competitively compensated to maintain quality.
@@ 3.2 Quality Control
Our quality control process primarily targets the automatically-crawled data, as the other two data collection strategies involve direct human involvement. To ensure accuracy, we recruited two professional annotators, each holding at least a bachelor’s degree and fluent in both Kazakh and Russian. They manually reviewed all questions to verify correctness and completeness. Any question containing errors or missing components (e.g., incomplete contexts or broken answer options) was discarded. Through this extensive human verification, every question included in KazMMLU undergoes manual validation, ensuring a high-quality dataset.
Additionally, we developed scripts to detect duplicates, verify metadata, and eliminate errors like duplicate questions, incorrect answer keys, and formatting issues, enhancing dataset reliability
@@ 3.3 Data Statistics
KazMMLU comprises 23,000 multiple-choice questions spanning two educational levels: high school and university. As shown in Figure 1, 48% of the dataset consists of high school-level questions in Kazakh, 36% in Russian, and 16% university-level questions in Russian. Table 2 outlines the subject distribution, covering STEM, Humanities, Social Sciences, and Languages. Notably, the Humanities, Social Sciences, and Language sections contain extensive Kazakhstan-specific knowledge. The dataset maintains a balanced distribution between Kazakh and Russian, reflecting the bilingual nature of Kazakhstan’s education system.
Table 3 presents the average question and answer lengths across educational levels and subject areas. While the overall question length remains relatively consistent between high school and university levels, answer lengths (in characters) tend to be longer at the university level. Additionally, questions in Humanities and STEM subjects are generally longer compared to those in Social Sciences and Languages.
@@ 4.1 Setup
We evaluate 27 multilingual LLMs of varying sizes in both zero-shot and few-shot settings. Our selection spans a range of architectures, including BLOOM Scao et al. (2023), Aya Üstün et al. (2024), Mistral Jiang et al. (2023), Gemma Team et al. (2024), Qwen Yang et al. (2025), GPT-4o OpenAI (2024), and DeepSeek V3 DeepSeek-AI (2024). To ensure regional relevance, we also include models tailored to specific languages: YandexGPT333yandex.cloud/en/services/yandexgpt and Vikhr Nikolich et al. (2024) for Russian, as well as KazakhLLM-8B444huggingface.co/issai/LLama-3.1-KazLLM-1.0-8B and Sherkala-Chat-8B Koto et al. (2025) for Kazakh. These regionally fine-tuned models are widely recognized for their focus on local languages.
For the evaluation purpose, we use two distinct prompt configurations to examine the effect of prompt language: (1) a Kazakh prompt with English (Latin script) alphabetic output and (2) an English prompt with English alphabetic output, as illustrated in Figure 4. For placeholders such as [Subject] and [Level], we use Kazakh in the Kazakh prompt and translate them into English for the English prompt. However, the question and answer choices remain in their original language (Kazakh or Russian). A complete example prompt is provided in Appendix B (Figure 10).
Following prior studies (Koto et al., 2023; Li et al., 2024), we adopt different answer selection methods based on model accessibility. For open-weight models, we apply the next-token prediction approach, computing probabilities for each multiple-choice option (A, B, C, D, or E) and selecting the one with the highest probability. This method is well-suited for autoregressive models that perform token-wise scoring. For closed-weight models (e.g., GPT–4o, Yandex-GPT, and DeepSeek V3), we use a free-text generation strategy, prompting the model to generate a textual response, from which the predicted answer is extracted via string matching. This approach is necessary due to the lack of direct token probability outputs in closed-source APIs. Regarding inference settings, we used the default configurations provided by their respective APIs or platforms. For evaluation, we use accuracy as the primary metric, following prior studies (Koto et al., 2023; Li et al., 2024).